# Inference-Time Text-to-Video Alignment with Diffusion Latent Beam Search

### v2

@@ Appendix E Theoretical Analysis on Lookahead Estimator
Appendix F List of Prompts
@@
Appendix G Detailed Analysis on Reward Function for Perceptual Video Quality
Figure 9 shows that different metrics in reward functions for perceptual video quality often exhibit negative or weak correlations. For example, dynamic degree tends to be negatively correlated with many other metrics, indicating that optimizing exclusively for one metric can either reduce motion dynamics or undermine temporal consistency and aesthetic quality. These findings underscore the need to balance potentially conflicting reward functions rather than prioritizing any single one in isolation, emphasizing the importance of a carefully calibrated approach to evaluating generated videos.
@@ Appendix G Detailed Analysis on Reward Function for Perceptual Video Quality
Appendix H Prompt of AI Feedback
@@
Appendix I Further Results for Calibrating Reward to Preference Feedback
I.1 Basic Prompts
Figure 12 and Figure 13 shows the two dimensional histogram and correlation between reward function and AI feedback from Gemini [14] and GPT-4o [13], and Figure 10 represent the coefficient of calibrated reward designed for GPT-4o. The relative weighting of the dynamic degree shifts in accordance with the prompt-specified dynamics, reflecting a pattern similar to that observed in the reward calibration approach with Gemini (DEVIL-high, DEVIL-medium, and DEVIL-static denote prompt sets characterized by high, medium, and low levels of dynamics, respectively). GPT-4o exhibits a stronger inclination toward dynamics than Gemini.
@@ I.1 Basic Prompts
I.2 Challenging Prompts
This section describes the reward calibration procedure and results for two challenging prompt sets, DEVIL-very-high and MovieGen, which were introduced to evaluate our method with larger T2V models. Following the methodology from for reward calibration with Latte (see Section 4), we generated 64 videos per prompt using Wan 2.1-1.3B [20]. Consistent with observations in Appendix I.1, using solely Text-Video Consistency is insufficient to fully capture AI feedback from Gemini (Figure 14). We choose the combination of weights $w_{i}$ to maximize correlation with Gemini’s evaluations. The coefficients of calibrated weights are shown in Figure 11.
@@ I.2 Challenging Prompts
Appendix J Correlation between VLM and Human Evaluation
As mentioned in the prior research [44, 45], VLMs such as Gemini and GPT-4o have been shown to exhibit a high correlation with human evaluations compared to other existing metrics. As an experiment, we measured the correlation between the AI feedback from these VLMs and human labels in the TVGE [72] dataset. As shown in Figure 15, Gemini achieved a correlation of 0.49, and GPT-4o achieved 0.51. Consequently, optimizing for these VLM rewards is a valid way to improve human-perceived quality, rather than merely “gaming” the metrics.
@@
Appendix K Qualitative Evaluation of Calibrated Reward
We provide example videos among individual rewards and VLM calibrated rewards in Figure 16. Videos selected solely on a single metric can over-optimize one aspect while neglecting others, whereas those chosen via VLM-calibrated rewards exhibit a more balanced quality. For instance, videos chosen solely based on temporal consistency (i.e., subject consistency and motion smoothness) or frame-by-frame quality (i.e., aesthetic quality, imaging quality) tend to lack dynamic movement, whereas those selected based on dynamic degree often lose temporal consistency. Evaluations relying on a single metric also fail to reflect the given prompt in some cases. Text-video consistency, which often exhibits a high correlation with VLM-based evaluation among individual metrics (Figure 3), is relatively effective in capturing the overall quality of a video. However, it may overlook certain aspects, such as frame-wise artifacts. In contrast, videos selected using VLM-calibrated rewards exhibit a more balanced overall quality.
@@
Appendix L Further Results for Diffusion Latent Beam Search
L.1 Scaling Search Budget with GPT-4o Calibrated Reward
We measure the performance in terms of a combinational reward calibrated to GPT-4o (Figure 17). DLBS improves all the calibrated rewards the best as the search budget $KB$ increases (especially $KB=16,32$), while BoN and GS, in some cases, eventually slow down or saturate the performance. Notably, an LA estimator with a small search budget ($KB=8,T^{\prime}=6$) is comparable to or even outperforms DLBS ($KB=32$).
@@
L.2 Scaling Search Budget to Larger Regimes
@@
L.3 Full Results for Scaling Trend of DLBS
@@
Figure 20: DLBS can improve the performance in any prompts or reward, as we increase the search budget $KB\in\{1,2,4,8,16,32\}$. In addition, we can see an optimal balance between the number of latent $K$ and the number of beam $B$ under the same budget. For instance, as we increase the budget to $KB=16$, we have a peak around $K=4$ or $K=8$, which is about 25–50% of the budget.
L.4 Further Analysis on Lookahead Estimator
@@
Figure 21 (Left) demonstrates that increasing the number of reward estimation steps $T^{\prime}$ in the LA estimator leads to improved reward prediction performance for $\mathbf{z}_{t}$ during the denoising process. This finding suggests that extending the LA steps enables a more effective search based on accurate reward predictions, particularly in the early stages of the denoising. As shown in Figure 21 (Right), enlarging the look-ahead horizon increases the reward gain to $T^{\prime}=6$; beyond this point, e.g., $T^{\prime}=20$ offers no significant benefit while multiplying the calculation. Accordingly, we fix $T^{\prime}=6$ in the main experiments, as it captures nearly all the attainable gains at minimal cost.
@@
Figure 21: (Left) Comparison of the reward estimation error for different LA steps $\mathbf{T^{\prime}}$. We evaluate the reward predicted by the LA estimator, which projects $\mathbf{z}_{t}$ to $\tilde{\mathbf{z}}_{0|\tilde{t}(0)}$ in $T^{\prime}$ steps (Algorithm 2) and computes $r^{\prime}(\tilde{\mathbf{z}}_{0|\tilde{t}(0)})$, against the actual reward obtained by projecting $\mathbf{z}_{t}$ to $\mathbf{z}_{0}$ in $t$ steps using a DDIM sampler ($\eta=0.0$) and evaluating $r^{\prime}(\mathbf{z}_{0})$. (Right) Impact of $T^{\prime}$ on search performance. Reward improves rapidly up to $T^{\prime}=6$ but saturates thereafter; using $T^{\prime}=20$ offers no measurable gain. These results show that a modest $T^{\prime}$ is sufficient in practice.
L.5 Ablation Study for Diffusion Steps
@@
Scaling Diffusion Steps  Figure 22 (Left) shows the performance when increasing the number of denoising steps $T$. Since DDIM exhibits fast convergence [42], BoN with larger $T$ does not improve the reward much. DLBS improves the performance when scaling denoising steps to $T=200$ more than BoN, which implies that DLBS benefits from larger computation and the other axes. However, as Figure 6 (Left) indicates, these gains are smaller than those obtained by widening the beam budget $KB$ or leveraging the LA estimator.
@@
Range of Diffusion Steps for Search  We investigate which range of diffusion steps DLBS should be applied to for effective search. Unlike Kim et al. [15], which apply GS only during the initial 5–10 steps, our results in Figure 22 (Right) show that applying DLBS throughout all steps leads to substantially better performance. This suggests that applying DLBS entirely is more effective than focusing on early stage.
@@
L.6 Ablation Study for $\eta$ in DDIM scheduler
@@
L.7 Alignment–Diversity Tradeoff
Alignment for diffusion models can steer desirable outputs, but it is said that the diversity of generated samples or the performance of original models often degrade [10, 37]. While inference-time search does not change or degrade the model itself, we here compare the diversity of samples among BoN, GS, and DLBS. We measure the sample diversity as the mean pairwise distance of ViCLIP [51] embeddings to quantify the diversity in videos, inspired by the approach for evaluating diversity in images [58]. Specifically, given $N$ generated video samples, we first extract ViCLIP embeddings $\mathbf{v}^{\text{video},(i)}$ for each sample $i$. The pairwise diversity score is then computed as the mean pairwise distance:
[…]
Figure 24 reveals that DLBS achieves high performance with higher diversity than BoN or GS. This exhibits a benefit from the wider possession and exploration of diffusion latents in DLBS.
Figure 24: Alignment–diversity tradeoff ($KB=32$), which is measured by the mean pairwise distance of ViCLIP embeddings. DLBS achieves high performance maintaining higher diversity.
L.8 Comparison with Other Inference-Time Search Methods
Concurrently with our work, inference-time search based on sequential Monte Carlo (SMC) has been proposed. We include a comparison with FK Steering [41] in Figure 25, where the resampling mechanism in the SMC-based methods does not occur frequently enough, preventing them from surpassing BoN performance, in our text-to-video experiments.
@@
L.9 DLBS with Larger Text-to-Video Models
We have tested our method on VideoCrafter2 [3] (1.9B) and CogVideoX-5B, 2B [19] and Wan 2.1-14B, 1.3B [20]. Our experiments confirm that the proposed approach yields significant improvements, indicating its effectiveness can be observed in larger video generation models.
@@
L.10 Further Results in AI and Human Evaluation
We show full results of evaluations using VideoScore [54], a metric trained on human judgments that evaluates videos at 8 fps across five dimensions and outputs scores ranging from 1.0 to 4.0 (see Figure 32). Under this metric, videos generated with the proposed search strategy consistently outperformed those generated without search.
Figure 32: Comparison of VideoScore [54]. We compare our approach in the $KB=1$, Best-of-N (BoN), and Diffusion Latent Beam Search (DLBS).
We also show full results of human judgment by three human evaluators (see Figure 33). These experiments confirmed that, regarding human preference (win rate), content generated with our search strategy consistently outperformed content produced without search.
Figure 33: Human evaluation results. We searched videos using the Gemini calibrated reward and asked three human evaluators to compare outputs from GS ($KB=1$) and DLBS-LA. "Win" indicates that the video generated by DLBS-LA was preferred.
L.11 Qualitative Results for DLBS
Qualitative results for DLBS are shown in https://sites.google.com/view/t2v-dlbs.
L.12 DLBS for Image Generation
We adopt PixArt-$\alpha$ [66] as our base text-to-image generation model. For evaluation, we directly reuse the prompt set of 45 common animal categories from prior works [73, 60] to ensure comparability. As a reward model, we employ the LAION aesthetic predictor [49] to assess image quality.
@@
Appendix M Extended Related Works

### v3

@@ Appendix F Theoretical Analysis on Lookahead Estimator
Appendix G List of Prompts
@@
Appendix H Detailed Analysis on Reward Function for Perceptual Video Quality
Figure 10 shows that different metrics in reward functions for perceptual video quality often exhibit negative or weak correlations. For example, dynamic degree tends to be negatively correlated with many other metrics, indicating that optimizing exclusively for one metric can either reduce motion dynamics or undermine temporal consistency and aesthetic quality. These findings underscore the need to balance potentially conflicting reward functions, rather than prioritizing any single one in isolation, and emphasize the importance of a carefully calibrated approach to evaluating generated videos.
@@ Appendix H Detailed Analysis on Reward Function for Perceptual Video Quality
Appendix I Prompt of AI Feedback
@@
Appendix J Further Results for Calibrating Reward to Preference Feedback
J.1 Basic Prompts
Figure 13 and Figure 14 show the two-dimensional histogram and correlation between reward function and AI feedback from Gemini [14] and GPT-4o [13], and Figure 11 represents the coefficient of calibrated reward designed for GPT-4o. The relative weighting assigned to the dynamic degree changes according to the dynamics grade of the prompt. Specifically, prompts with a high dynamics grade, i.e., DEVIL-high, place greater weight on the dynamic degree. In contrast, prompts that describe slight motion, i.e., DEVIL-medium and DEVIL-static, place a smaller weight on it. This behavior matches the pattern observed in reward calibration with Gemini (Figure 4). GPT-4o exhibits a stronger inclination toward dynamics than Gemini.
@@
J.2 Challenging Prompts
This section describes the reward calibration procedure and results for two challenging prompt sets, DEVIL-very-high and MovieGen, which were introduced to evaluate our method with larger T2V models, such as CogVideoX [19] and Wan 2.1 [20]. Following the methodology for reward calibration with Latte (see Section 4), we generated 64 videos per prompt using Wan 2.1-1.3B [20]. Consistent with observations in Appendix J.1, using solely text-video consistency is insufficient to fully capture AI feedback from Gemini (Figure 15). We choose the combination of weights $w_{i}$ to maximize correlation with Gemini’s evaluations. The coefficients of calibrated weights are shown in Figure 12.
@@
J.3 Cost of Reward Calibration
As described in Section 4.1 and Appendix J.2, we generate 64 videos per prompt using pre-trained Latte and Wan 2.1 models. Compared to naively querying VLMs at every inference step, our calibration approach is substantially more cost-efficient, since the VLM queries are amortized through a one-time weight estimation. Table 2 summarizes the difference in per-prompt query count and execution time when applying DLBS (KB=32K\!B=32). These results demonstrate that reward calibration reduces the number of VLM queries, making large-scale search with DLBS computationally feasible.
Table 2: Comparison of query count and execution time between naive VLM queries during search and reward calibration. Assuming 15 seconds per VLM query.
Method Query Count Exec. Time (sec)
Querying VLMs during Search (KB=32K\!B=32) $T{=}50\times K\!B{=}32=1600$ $\approx 102{,}400$
Reward Calibration $\bf{64}$ $\bf{\approx 960}$
J.4 Generalization of Reward Calibration across prompts
Video generation inherently involves trade-offs between fundamental properties such as dynamics and consistency (Appendix H), which may require category-specific calibration for optimal performance. However, despite these domain-specific requirements, we hypothesize that calibrated rewards can generalize to some extent across different datasets, as they are based on shared principles of perceptual quality. To test the out-of-domain transferability, we conducted additional experiments applying the reward weights calibrated on DEVIL-high and DEVIL-medium to MSRVTT-test prompts (Table 3). We used Latte [18] as a base model and evaluated the results using VideoScore [55], a human preference-trained evaluator, measuring five key metrics along with their corresponding average scores. With the DEVIL-high reward, we can enhance other metrics while maintaining dynamics. DEVIL-medium reward, which is a closer domain to MSRVTT-test, shows a different trade-off pattern. While it slightly reduces dynamics, it significantly improves other metrics and achieves a higher average score than the MSRVTT-test reward, demonstrating higher transferability.
Table 3: Out-of-domain prompt generalization. Rewards calibrated on DEVIL-high/medium applied to MSRVTT-test prompts. All metrics are derived from VideoScore [55]. VQ = Visual Quality; TC = Temporal Consistency; $D$ = Dynamic Degree; T2V Align. = Text-to-video Alignment; FC = Factual Consistency.
(R: Reward, P: Prompt) VQ TC $D$ T2V Align. FC Average
Latte 2.32 2.01 2.91 2.67 2.07 2.40
+ DLBS (KB=8K\!B=8)
with R=MSRVTT, P=MSRVTT 2.50 2.27 2.88 2.74 2.28 2.53
with R=DEVIL-medium, P=MSRVTT 2.49 2.26 2.89 2.73 2.31 2.54
with R=DEVIL-high, P=MSRVTT 2.36 2.03 2.90 2.68 2.10 2.42
Appendix K Correlation between VLM and Human Evaluation
As mentioned in prior research [44, 45, 46], evaluation from VLMs such as Gemini and GPT-4o exhibits a high correlation with human assessment compared to other existing metrics. As an experiment, we measured the correlation between the AI feedback from these VLMs and human labels in the TVGE dataset [74]. As shown in Figure 16, Gemini achieved a correlation of 0.49, and GPT-4o achieved 0.51. Consequently, optimizing for these VLM rewards is a valid way to improve human-perceived quality, rather than merely “gaming” the metrics.
@@
For a deeper analysis of failure cases, we qualitatively examined the top 5% outliers between human preference labels in the TVGE dataset and VLM (Gemini) evaluation (Figure 17). As far as we observed, VLM sometimes makes subtle mistakes, but we did not see any critical failures.
Figure 17: Misaligned cases of Gemini-based evaluation with human preferences. (Top) For prompts specifying text, such as "the words ’KEEP OFF THE GRASS’ on a sign next to a lawn," the VLM was significantly harsher than human evaluation on text rendering (VLM: 2/10, Human: 4.5/5). (Middle) For prompts specifying quantities, such as "Four friends have a picnic, enjoying six sandwiches and two bottles of juice," the VLM was more lenient than human evaluation (VLM: 8/10, Human: 1.5/5). (Bottom) For "fashion portrait shoot of a girl in colorful glasses, a breeze moves her hair," despite missing arms in the generated person, the VLM was misled by distracting background patterns, possibly mistaking them for curtain-like elements that obscure the arms behind the background (VLM: 8/10, Human: 1.2/5).
Appendix L Qualitative Evaluation of Calibrated Reward
We provide best-of-64 videos by individual rewards and VLM calibrated rewards in Figure 18. Videos selected solely on a single metric can over-optimize one aspect while neglecting others, whereas those chosen via VLM-calibrated rewards exhibit a more balanced quality. For instance, videos chosen solely based on temporal consistency (i.e., subject consistency and motion smoothness) or frame-by-frame quality (i.e., aesthetic quality, imaging quality) tend to lack dynamic movement, whereas those selected based on dynamic degree often lose temporal consistency. Evaluations relying on a single metric also fail to reflect the given prompt in some cases. Text-video consistency, which often exhibits a high correlation with VLM-based evaluation among individual metrics (Figure 3), is relatively effective in capturing the overall quality of a video. However, it may overlook certain aspects, such as frame-wise artifacts. In contrast, videos selected using VLM-calibrated rewards exhibit a more balanced overall quality.
@@
Appendix M Further Results for Diffusion Latent Beam Search
M.1 Scaling Search Budget with GPT-4o Calibrated Reward
We measure the performance using a reward calibrated to GPT-4o (Figure 19). DLBS improves all the calibrated rewards the best as the search budget $KB$ increases (especially $KB=16,32$), while BoN and GS, in some cases, eventually slow down or saturate the performance. Notably, an LA estimator with a small search budget ($KB=8,T^{\prime}=6$) is comparable to or even outperforms DLBS ($KB=32$).
@@
M.2 Scaling Search Budget to Larger Regimes
@@
M.3 Full Results for Scaling Trend of DLBS
@@
Figure 22: DLBS can improve the performance in any prompts or reward, as we increase the search budget $KB\in\{1,2,4,8,16,32\}$. In addition, we can see an optimal balance between the number of latent $K$ and the number of beam $B$ under the same budget. For instance, as we increase the budget to $KB=16,32$, we have a peak around $K=4,8,16$, which is about 25–50% of the budget.
M.4 Further Analysis on Lookahead Estimator
@@
Figure 23 (Left) demonstrates that increasing the number of reward estimation steps $T^{\prime}$ in the LA estimator leads to improved reward prediction performance for $\mathbf{z}_{t}$ during the denoising process. This finding suggests that extending the LA steps enables a more effective search based on accurate reward predictions, particularly in the early stages of the denoising. As shown in Figure 23 (Right), enlarging the look-ahead horizon increases the reward gain to $T^{\prime}=6$; beyond this point, e.g., $T^{\prime}=20$ offers no significant benefit while multiplying the computational cost. Accordingly, we fix $T^{\prime}=6$ in the main experiments, as it captures nearly all the attainable gains at minimal cost. These results were obtained on Latte [18] using a DDIM sampler.
@@
Figure 23: (Left) Comparison of the reward estimation error for different LA steps $T^{\prime}$. We evaluate the reward predicted by the LA estimator, which projects $\mathbf{z}_{t}$ to $\tilde{\mathbf{z}}_{0|\tilde{t}(0)}$ in $T^{\prime}$ steps (Algorithm 2) and computes $r^{\prime}(\tilde{\mathbf{z}}_{0|\tilde{t}(0)})$, against the actual reward obtained by projecting $\mathbf{z}_{t}$ to $\mathbf{z}_{0}$ in $t$ steps using a DDIM sampler ($\eta=1.0$) and evaluating $r^{\prime}(\mathbf{z}_{0})$. (Right) Impact of $T^{\prime}$ on search performance. Reward improves rapidly up to $T^{\prime}=6$ but saturates thereafter; using $T^{\prime}=20$ offers no measurable gain. These results show that a modest $T^{\prime}$ is sufficient in practice.
To verify that this behavior is not specific to DDIM sampler, we conducted the same ablation with an SDE-DPMSolver++ [56] on Wan 2.1-1.3B [20] (Figure 24). Note that the notation follows Algorithm 4. Specifically, $M^{\prime}$ in the SDE-DPMSolver++ setting corresponds to the $T^{\prime}$ used with the DDIM sampler. We observed the same pattern shown in Figure 23. Increasing the look-ahead horizon $M^{\prime}$ monotonically improves the LA estimator’s reward prediction. Search reward gain up to roughly $M^{\prime}=6$, after which gains saturate. For example, $M^{\prime}=12$ yields no measurable improvement while incurring substantially higher cost. This cross-sampler and cross-model consistency provides a practical guideline for choosing $M^{\prime}$: a modest horizon ($\approx 6$) captures nearly all attainable benefit.
Figure 24: (Left) Comparison of the reward estimation error for different LA steps $M^{\prime}$. We evaluate the reward predicted by the LA estimator, which projects $\mathbf{z}_{t_{s}}$ to $\tilde{\mathbf{z}}_{{t_{M}}|t_{\tilde{s}(0)}}$ in $M^{\prime}$ steps (Algorithm 4) and computes $r^{\prime}(\tilde{\mathbf{z}}_{{t_{M}}|t_{\tilde{s}(0)}})$, against the actual reward obtained by projecting $\mathbf{z}_{t_{s}}$ to $\mathbf{z}_{t_{M}}$ in (Ms)(M-s) steps using a SDE-DPMSolver++ [56] and evaluating $r^{\prime}(\mathbf{z}_{M})$. (Right) Ablation study of $M^{\prime}$ on search performance with SDE-DPMSolver++ [56] on Wan 2.1-1.3B [20]. Reward improves rapidly up to $M^{\prime}=6$ but saturates thereafter; using $M^{\prime}=12$ offers no measurable gain. These results show that a modest $M^{\prime}$ is sufficient in practice.
M.5 Ablation Study for Diffusion Steps
@@
Scaling Diffusion Steps  Figure 25 (Left) shows the performance when increasing the number of denoising steps $T$. Since DDIM exhibits fast convergence [42], BoN with a larger $T$ does not improve the reward much. DLBS improves performance when scaling denoising steps to $T=200$ more than BoN, which implies that DLBS benefits from larger computational resources in denoising. However, as Figure 6 (Left) indicates, these gains are smaller than those obtained by widening the beam budget $KB$ or leveraging the LA estimator.
@@
Range of Diffusion Steps for Search  We investigate which range of diffusion steps DLBS should be applied to for effective search. Unlike Kim et al. [15], which applies GS only during the initial 5–10 steps, our results in Figure 25 (Right) show that applying DLBS throughout all steps leads to substantially better performance. This suggests that applying DLBS entirely is more effective than focusing on the early stage.
@@
M.6 Ablation Study for $\eta$ in DDIM scheduler
@@
M.7 Comparison with Gradient-Based Search Methods
Applicability to diverse reward models  One advantage of our zero-order search framework is its applicability to reward models for which computing gradients is computationally prohibitive (e.g., large-scale VLMs) or fundamentally impossible (e.g., human evaluators or external API-based models). To demonstrate this, we simulated search with human feedback by employing VideoScore [55], a VLM-based reward model trained on human evaluations, as the reward function. As shown in Table 4, DLBS-LA with VideoScore as the evaluator achieved substantial improvements over the vanilla baseline, suggesting that high-performance VLMs or human evaluators can, in principle, be directly incorporated as reward functions in our framework.
Table 4: Results of DLBS-LA with VideoScore as the evaluator, illustrating its applicability to reward models for which computing gradients is computationally prohibitive (e.g., large-scale VLMs) or fundamentally impossible (e.g., human evaluators or external API-based models).
Method VideoScore
Latte 2.40
+ DLBS-LA (KB=4K\!B=4, T=6T’=6) with VideoScore 2.69
Efficiency in time and memory  We further compared DLBS with DAS [61], a first-order gradient-based method based on Sequential Monte Carlo. Experiments were conducted on Stable Diffusion 1.5 [21] with LAION Aesthetic V2 [50] as the reward model, using an NVIDIA RTX 6000 Ada (48GB). Table 5 shows that under the assumption of equal execution time, DLBS (a zero-order method) takes the lead because it can have a larger search budget, which refers to the number of particles used for search, i.e., KBK\!B for DLBS and $N$ for DAS. Our observation that a zero-order method achieves better performance than a first-order method under the equal execution time aligns with prior findings on inference-time search for image generation [41].
Gradient-based search methods also exhibit significant increases in memory usage due to gradient computations required for the reward function and the VAE decoder. In other words, gradient-based methods are actually inefficient in terms of memory cost. The results shown in Table 5 are based on a single evaluator and a single frame (i.e., image generation). Note that for video generation, as mentioned in Figure 3, a single evaluator metric does not correlate well with perceptual quality, necessitating the combination of multiple evaluators, which roughly multiplies the memory requirements for gradient calculation. Additionally, video generation models do not decode just one frame from the VAE (maximum frames are 81 for Wan 2.1 [20] and 49 for CogVideoX [19]). The gradient increase would be significantly larger in video generation than in image generation, making gradient-based methods almost impossible in practice. For reference, the memory usage of vanilla Latte, DLBS-LA, and DAS in Table 6.
Table 5: Comparison between DLBS and DAS on Stable Diffusion 1.5 with LAION Aesthetic V2 reward. DLBS (a zero-order method) achieves better performance than DAS (a first-order method) under equal execution time.
Method Score Time (sec) Memory (GB)
SD 1.5 5.81 2 4.3
+ DAS ($N=8$) 6.59 108 14.2
+ DLBS ($K=8,B=2$) 6.63 103 5.9
+ DAS ($N=16$) 6.68 220 14.2
+ DLBS ($K=16,B=2$) 6.69 209 5.9
Table 6: Memory usage of search methods on Latte. The memory advantage of DLBS-LA (a zero-order method) becomes more critical in video generation.
Method Latte + DLBS-LA + DAS
Memory Usage (GB) 16.3 38.9 >>48.0 (OOM)
M.8 Comparison with Other Inference-Time Search Methods
Concurrently with our work, inference-time search based on zero-order sequential Monte Carlo (SMC) has been proposed. We include a comparison with FK Steering [41] in Figure 27, where the resampling mechanism in the SMC-based methods does not occur frequently enough, preventing them from surpassing BoN performance, in our text-to-video experiments.
@@
M.9 Scalability to Long Videos
To demonstrate that our method is scalable even when extending the frame count to the model’s maximum, we conducted experiments with the maximum frames for CogVideoX-5B (49 frames, 6 seconds) and Wan 2.1-1.3B (81 frames, 5 seconds). The reward, which was calibrated using 33-frame, 2-second videos generated by Wan 2.1-1.3B, was applied as-is. As shown in Table 7 and Table 8, we confirmed that even with longer frames, the reward values could be improved more efficiently than the BoN baseline.
Table 7: Scalability results on CogVideoX-5B with 49 frames, 6 seconds using DEVIL-very-high prompts.
Method KBKB Reward Inference Compute (NFE)
CogVideoX-5B 1 0.429 50
+ BoN 64 0.474 3200
128 0.478 6400
+ DLBS 16 0.481 1200
32 0.490 2400
+ DLBS-LA 8 0.497 2500
16 0.517 4900
Table 8: Scalability results on Wan 2.1-1.3B with 81 frames, 5 seconds using MovieGen prompts.
Method KBKB Reward Inference Compute (NFE)
Wan 2.1-1.3B 1 0.313 50
+ BoN 128 0.357 6400
256 0.360 12800
+ DLBS 16 0.357 1200
32 0.371 2400
+ DLBS-LA 8 0.373 2500
16 0.393 5000
M.10 DLBS with Larger Text-to-Video Models
We have tested our method on VideoCrafter2 [3] (1.9B) and CogVideoX-5B, 2B [19] and Wan 2.1-14B, 1.3B [20]. Our experiments confirm that our DLBS and DLBS-LA yield significant improvements, indicating their effectiveness can be observed in larger video generation models.
@@
M.11 Further Results in AI and Human Evaluation
We show full results of evaluations using VideoScore [55], a metric trained on human judgments that evaluates videos at 8 fps across five dimensions and outputs scores ranging from 1.0 to 4.0 (see Figure 34). Under this metric, videos generated with our DLBS consistently outperformed those generated without search and with BoN.
Figure 34: DLBS on calibrated reward also improves another qualitative metric, the most, VideoScore [55], which is not involved in a reward calibration.
We also show additional results of human judgment by three human evaluators (see Figure 35). These experiments confirmed that, regarding human preference (win rate), content generated with our search strategy consistently outperformed content produced without search.
Figure 35: Human evaluation results. We searched videos using the Gemini calibrated reward and asked three human evaluators to compare outputs from GS ($KB=1$) and DLBS-LA ($KB=8$, T=6T’=6). "Win" indicates that the video generated by DLBS-LA was preferred.
M.12 Qualitative Results for DLBS
Qualitative results are shown in https://sites.google.com/view/t2v-dlbs.
M.13 DLBS for Image Generation
We adopt PixArt-$\alpha$ [69] as our base text-to-image generation model. For evaluation, we directly reuse the prompt set of 45 common animal categories from prior works [75, 63]. As a reward model, we employ the LAION aesthetic predictor [50] to assess image quality.
@@
Appendix N Extended Related Works